A particular organic molecule forms an electric dipole by possessing an effective charge separation of 0.193 nm for a pair of +1.60 x 1019-C charges. What is the magnitude of the molecules electric dipole moment? Number C m Find the magnitude of the torque that acts on the molecule when it is immersed in a uniform electric field of 3.77 x105 N/C with its electric dipole vector at an angle of 58.5° from the direction of the field. Number N. m Consider all possible angles between the electric field and the electric dipole vector. What is the greatest magnitude of torque that the electric field can exert on this molecule? Number N- m
Answer #1

Solution 1)

Magnitude of the electric dipole moment

P= q*d

Where, q= magnitude of charge

D= distance between them

So, p= (1.6*10^-19 C)*(0.193*10^-9 m)

P= 3.088 * 10^-29 C-m (Ans)